OpenStudy (anonymous):

<1 and <2 are supplementary angles. m<1 = 4y + 7, and m<2 = 9y + 4. Find m<2. I figured out that y = 13 so that would make m<2 = 121, but I need the work shown and I don't know how to do the work.

OpenStudy (anonymous):

@DuckDynastyfan923

OpenStudy (godlovesme):

do u have a figure? @AquaStone316

OpenStudy (anonymous):

No @Godlovesme

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

Great

OpenStudy (anonymous):

So the two angles are supplementary which means that the add up to 180

OpenStudy (anonymous):

Yeah, I already figured out that part and that y = 13 and that would make m<1 59 and m<2 121. But I need the work shown in order to answer because I sometimes just figure it without the work but the school wants me to do the work and not just the answer.

OpenStudy (anonymous):

Yeah I understand that I was gonna help you with that, because I had that problem too

OpenStudy (anonymous):

So with that you should show how you got what y equals then you show the answer of both angles when you plug the answer for y into the equations

OpenStudy (anonymous):

That's what I have trouble with, I forget how to show the work for how I got what y equals to.

OpenStudy (anonymous):

@WOLFTHORN

OpenStudy (anonymous):

i will show you what it looks like ok

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

(4y+7)+(9y+4)=180 13y+11=180 13y=169 y=13 <1= 4(13)+7=59 <2=9(13)+4=121 and that is how you show your work

OpenStudy (anonymous):

Ok thank you.